# Deep Reinforcement Learning for Robotic Bipedal Locomotion: A Brief Survey

### v6

@@ I Introduction
As a result, bipedal robots hold significant potential for real-world applications [2024_tong_advancements_review]. In manufacturing, they can perform tasks efficiently without requiring additional tools, thereby enhancing productivity and reducing labour demands [industrial_robotics_applications, collaborative_industrial, 2023_vidoeo_demo_digit_locomaniputlation]. Their agility is particularly advantageous in complex environments such as multi-level workplaces. Bipedal robots are also well suited to tasks that involve the use of human-designed tools, making them valuable for assisting in daily activities, healthcare, and rehabilitation [2019_bingjing_human_exoskeleton]. Moreover, they show considerable promise in search-and-rescue operations, where they can navigate hazardous and unpredictable terrains [2015_underwaterrobot_review, quadSpace1, quadSpace2].
Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [2017_biped_model-based_method_review, 2021_Jenna_bipedrobot_review, 2021_Justin_review_recentporgressinleggedrobot]. Early methods, such as the Linear Inverted Pendulum Model (LIPM) [2013_3DSLIPmodel_running], provided simplified representations of the dynamics involved in bipedal motion, enabling easier analysis and control. As research progressed, full dynamic models were introduced to better capture the complexities of real-world locomotion. Advanced methods such as Model Predictive Control (MPC) [2022_Hou_drlasparametertoMPC, 2023_junheng_steppingstones_MPCWBC] and Trajectory Optimisation (TO) [2016_LIPM_trajectoryoptimization, 2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2017_biped_model-based_method_review] exploit predefined dynamic models to solve constrained optimal-control problems that plan footsteps, centre-of-mass (CoM) motion, and contact forces. While model-based approaches offer rapid convergence and predictive capabilities, they often struggle in dynamically complex and uncertain environments where adaptability is essential.
Reinforcement learning (RL)-based methods, particularly deep reinforcement learning (DRL), are effective in optimising robot control policies through direct interaction with the environment [2020_systematic_review], which provides a distinct advantage. Unlike model-based approaches, which rely on predefined dynamics and may fail under unforeseen conditions, DRL enables robots to autonomously discover control strategies through trial and error, achieving greater adaptability and robustness in diverse environments. In addition, hybrid methods that combine model-based and learning-based techniques further enhance planning and control by leveraging the strengths of both paradigms.
@@
The evolution of RL in bipedal robotics has largely advanced through the end-to-end learning paradigm. Early studies in 2004 applied simple policy-gradient methods to 2D bipeds [2004_Tedrake_RL_bipedal_policygradient, 2004_model-based_RL_learnpoincaremap_simple_bipedrobot], while later breakthroughs in DRL enabled policy training in high-fidelity physics simulators [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2018_wenhao_yu_DRL_withoutpredefine_symmetrygait_PPO_jointanglePD_]. As robotic hardware matured, an increasing variety of bipedal and humanoid platforms emerged, supporting extensive evaluation of DRL-based locomotion across diverse morphologies, as illustrated in Fig. 1. This evolution marked the transition from purely simulation-based training to simulation-to-real (sim-to-real) transfer, where policies trained in simulators are deployed on physical robots. In 2020, the first successful sim-to-real transfer of an end-to-end DRL locomotion policy was achieved on the 3D torque-controlled bipedal robot Cassie [2020_Xie_firstsim2real_]. Subsequent work explored two principal learning paradigms: reference-based learning, which leverages TO-generated data or motion-capture data to guide policy training [2021_michael_velocity-basedcontrol_motioncapturedata, 2024_speratebody_expressive_allmotion, 2024_tang_humanmimic_adversarial, 2024_Zhang_wholebody_adversarial_motion_priors]; and reference-free learning, where policies are trained entirely from scratch to autonomously discover control strategies [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie]. These developments demonstrate that end-to-end frameworks can achieve robust and versatile locomotion skills across complex terrains and dynamic environments [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2023_Chongben_jumping_DDPG_multi-task_network_simulationonly, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining].
@@
Similarly, hierarchical structures have garnered significant interest. Within this subset, the hybrid approach combines RL-based and model-based methods to enhance both planning and control strategies. Hybrid architectures often integrate learning-based and model-based modules to combine adaptability with physical consistency. One representative design couples a learned HL planner with an LL model-based controller, forming a cascade-structure or deep-planning hybrid scheme [2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2021_duan_DRL_task-spaceaction_hiarachycotnrolscheme_inversedynamiccontroller, 2022_DRL_cascade_motionplanningpolicy_HZD_torso_PPOwithRNN]. Alternatively, DRL feedback-control hybrids embed learned control policies within model-based feedback loops to enhance tracking precision and disturbance rejection [2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning]. Learned hierarchical control schemes [2017_DRL_survey] decompose locomotion into multiple layers, each focusing on specific functions such as navigation and fundamental locomotion skills [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2023_Wei_learned_hierarchy_framework_wheeled_bipedalrobot]. To provide a clearer overview of the current landscape, we categorise existing DRL frameworks as shown in Fig. 2.
@@
Current progress across both end-to-end and hierarchical paradigms indicates that a unified framework for DRL-based bipedal locomotion is still far from being realised. Establishing such a framework is essential for consolidating diverse learning pipelines, standardising evaluation metrics, and enabling transferable locomotion capabilities across different robot morphologies. As locomotion tasks become increasingly complex, ranging from basic stabilisation to dynamic parkour and loco-manipulation, the need for consistent benchmarking has intensified.The DARPA Robotics Challenge exemplified this trend by introducing one of the first large-scale evaluation platforms for bipedal humanoids performing real-world locomotion and manipulation tasks [2015_DAPRA_Challenge_1], highlighting the importance of robustness and practical deployment.
Although several reviews discuss RL for general robotics [2020_systematic_review] and model-based methods for bipedal robots [2017_biped_model-based_method_review, 2021_Jenna_bipedrobot_review, 2021_Justin_review_recentporgressinleggedrobot], none specifically focus on DRL-based frameworks for bipeds. To address this gap, this survey reviews relevant literature according to the following selection criteria: (1) studies that investigate DRL frameworks specifically for bipedal robots; (2) research involving both simulated and physical bipedal robots; (3) approaches that improve policy transfer from simulation to real-world environments; and (4) publications from the last five years (2018–April 2024) sourced from reputable databases and conferences, including Google Scholar, IEEE Xplore, Web of Science, arXiv, and major robotics venues such as CoRL, RSS, ICRA, IROS, and Humanoids.
@@ II End-to-end framework
The end-to-end DRL framework represents a holistic approach in which a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly maps sensory inputs $\mathcal{X}$, such as images, LiDAR data, or proprioceptive feedback [2017_compare_actionspace_RL], together with user commands [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie] or pre-defined references [2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter], into joint-level control actions $\mathcal{U}$. Here, $\mathcal{X}$ represents the sensory input space, $\mathcal{U}$ refers to the space of control actions, and $\pi(\cdot)$ denotes the policy function. This framework obviates the need for manually decomposing the problem into sub-tasks, streamlining the control process.
@@
Methods
Works
Capabilities
Characteristic
Advantages and Disadvantages
Implementation Flow Chart
Residual learning
[2018_xie_OSUDRL_feedback_DRL_reference-based]
[2020_Xie_firstsim2real_]
Forward walk
unidirectional walk
Adding a residual term to the known motor positions at the current time step.
A: Fast convergence speed D: Requires high-quality predefined reference, limits to specific motions, and lacks robustness to complex terrains.
Guided learning
[2020_Siekmann_drl_e2e_s2r_cassie]
[2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter]
[2023_UCB_Cassie_RL_Jumping]
[2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining]
Forward walk
Versatile walk
Versatile jump
Versatile motions
Mimic the predefined reference and directly specifies joint-level commands.
A: Accelerates the learning process and is robust across terrains. D: Limits to the predefined motions and lacks adaptability to unforeseen changes in environment.
Reference-free learning
[2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie]
[2021_siekmann_blind_DRL_stair]
[2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark]
Periodic motions
Stepping stones walk
Visual walk
Learn locomotion skills from scratch without any prior knowledge.
A: High potential for gait exploration, highly robust to complex terrain D: Requires intensive reward shaping for gait patterns and is relatively expensive in computational resources
@@ II-A1 Residual learning
Introduced in 2018, a residual learning framework for the bipedal robot Cassie marked a significant advancement [2018_xie_OSUDRL_feedback_DRL_reference-based]. This framework allowed the robot to walk forward by incorporating a policy trained via Proximal Policy Optimisation (PPO) algorithms, as detailed in Appendix A. The policy receives the robot’s states and reference inputs, outputting a residual term that augments the reference at the current timestep. These modified references are then processed by a PD controller to set the desired joint positions. Although this framework has improved the robot’s ability to perform tasks beyond standing [2018_yang_push_recover_wholebody], it has yet to be physically deployed on a bipedal robot. As a result, it remains impractical for managing walking at varying speeds and is constrained to movement in a single direction.
To transition this framework to a real robot, a sim-to-real strategy based on the previous model was demonstrated, where the policy, trained through a residual learning approach, was subsequently applied to a physical bipedal robot [2020_Xie_firstsim2real_]. Compared to model-based methods, this training policy achieves faster running speeds on the same platform, underlining the considerable potential of DRL-based frameworks. However, the robot’s movements remain constrained to merely walking forward or backward.
A unique residual learning approach was introduced to enable omnidirectional walking, where the policy adds a residual term to the current joint positions, allowing gradual omnidirectional walking [2021_Diego_deepwalk_curriculum_learning_sim2real]. In this case, the desired reference is the robot’s current joint positions, which makes the approach distinctive. However, this also limits the policy’s ability to explore more diverse motions, restricting it to a single slow walking pattern.
Residual learning enhances an existing control policy by taking current joint positions or states and applying a residual action to adjust reference actions for better performance. Compared to other learning approaches that directly output joint positions, it is highly sample efficient [2021_duan_DRL_task-spaceaction_hiarachycotnrolscheme_inversedynamiccontroller]. However, when predefined references are unstable or of low quality, residual learning may struggle, especially on complex terrains, as the action space is bounded by the reference, limiting the ability to handle unpredictable or uneven terrains.
@@ II-A2 Guided learning
A sim-to-real framework that employs periodic references to initiate the training phase was proposed in [2020_Siekmann_drl_e2e_s2r_cassie]. In this framework, the action space directly maps to the joint angles, and desired joint positions are managed by joint PD controllers. The framework also incorporates a Long Short-Term Memory (LSTM) network, as detailed in the Appendix A, which is synchronised with periodic time inputs. However, this model is limited to a single locomotion goal: forward walking. A more diverse and robust walking DRL framework that includes a Hybrid Zero Dynamics (HZD) gait library was demonstrated [2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter], achieving a significant advancement by enabling a single end-to-end policy to facilitate walking, turning, and squatting.
@@
Despite these advancements, the parameterisation of reference motions introduces constraints that limit the flexibility of the learning process and the policy’s response to disturbances. To broaden the capabilities of guided learning policies, a framework capable of handling multiple targets, including jumping, was developed [2023_UCB_Cassie_RL_Jumping]. This approach introduced a novel policy structure that integrates long-term input/output (I/O) encoding, complemented by a multi-stage training methodology that enables the execution of complex jumping manoeuvrers. An adversarial motion priors approach, employing a style reward mechanism, was also introduced to facilitate the acquisition of user-specified gait behaviours [2024_Zhang_wholebody_adversarial_motion_priors]. This method improves the training of high-dimensional simulated agents by replacing complex hand-designed reward functions with more intuitive controls.
@@
While previous works primarily focused on specific locomotion skills, a unified framework that accommodates both periodic and non-periodic motions was further developed [2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining] based on the foundational work in [2023_UCB_Cassie_RL_Jumping]. This framework enhances the learning process by incorporating a wide range of locomotion skills and introducing a dual I/O history approach, marking a significant breakthrough in creating a robust, versatile, and dynamic end-to-end framework. However, experimental results indicate that the precision of locomotion features, such as velocity tracking, remains suboptimal.
@@
Guided learning methods expedite the learning process by leveraging expert knowledge and demonstrating the capacity to achieve versatile and robust locomotion skills. Through the comprehensive evaluation [2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining], it is demonstrated that guided learning employs references without complete dependence on them. Conversely, residual learning exhibits failures or severe deviations when predicated on references of inferior quality. This shortfall stems from the framework’s dependency on adhering closely to the provided references, which narrows its learning capabilities.
@@
However, the benefits of reference-based learning come with inherent limitations. Reliance on predefined trajectories often confines the policy to specific gaits, restricting its capacity to explore a broader range of motion possibilities [2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter, 2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction]. Moreover, such methods exhibit reduced adaptability when confronted with novel environments or unforeseen perturbations. These limitations are further compounded by the difficulty of acquiring high-quality and task-relevant demonstrations.
@@
Common sources of prior knowledge include TO [2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real, 2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining], human motion capture [2024_speratebody_expressive_allmotion], teleoperation [2023_seo_deep_loco_manipulation, 2024_stanford_humanplus_hierarchical_humanoidshadowingimitation], and scripted controllers [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task]. While informative, these demonstrations often require adaptation due to embodiment mismatch or limited generalisability. Motion retargeting [2019_iit_wholebody_retargeting, 2024_tang_humanmimic_adversarial, 2017_Ayusawa_motionretarget_optimization], as one of promising direction, addresses this by converting human-centric motions into robot-feasible trajectories, yet it still struggles with preserving fidelity and adapting across morphologies.

### v7

@@ I Introduction
As a result, bipedal robots hold significant potential for real-world applications [110]. In manufacturing, they can perform tasks efficiently without requiring additional tools, thereby enhancing productivity and reducing labour demands [26, 120, 1]. Their agility is particularly advantageous in complex environments such as multi-level workplaces. Bipedal robots are also well suited to tasks that involve the use of human-designed tools, making them valuable for assisting in daily activities, healthcare, and rehabilitation [9]. Moreover, they show considerable promise in search-and-rescue operations, where they can navigate hazardous and unpredictable terrains [11, 94, 87].
Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [37, 91, 14]. Early methods, such as the Linear Inverted Pendulum Model (LIPM) [113], provided simplified representations of the dynamics involved in bipedal motion, enabling easier analysis and control. As research progressed, full dynamic models were introduced to better capture the complexities of real-world locomotion. Advanced methods such as Model Predictive Control (MPC) [42, 59] and Trajectory Optimisation (TO) [40, 60, 37] exploit predefined dynamic models to solve constrained optimal-control problems that plan footsteps, centre-of-mass (CoM) motion, and contact forces. While model-based approaches offer rapid convergence and predictive capabilities, they often struggle in dynamically complex and uncertain environments where adaptability is essential.
Reinforcement learning (RL)-based methods, particularly deep reinforcement learning (DRL), are effective in optimising robot control policies through direct interaction with the environment [54], which provides a distinct advantage. Unlike model-based approaches, which rely on predefined dynamics and may fail under unforeseen conditions, DRL enables robots to autonomously discover control strategies through trial and error, achieving greater adaptability and robustness in diverse environments. In addition, hybrid methods that combine model-based and learning-based techniques further enhance planning and control by leveraging the strengths of both paradigms.
@@
The evolution of RL in bipedal robotics has largely advanced through the end-to-end learning paradigm. Early studies in 2004 applied simple policy-gradient methods to 2D bipeds [108, 78], while later breakthroughs in DRL enabled policy training in high-fidelity physics simulators [84, 83, 125]. As robotic hardware matured, an increasing variety of bipedal and humanoid platforms emerged, supporting extensive evaluation of DRL-based locomotion across diverse morphologies, as illustrated in Fig. 1. This evolution marked the transition from purely simulation-based training to simulation-to-real (sim-to-real) transfer, where policies trained in simulators are deployed on physical robots. In 2020, the first successful sim-to-real transfer of an end-to-end DRL locomotion policy was achieved on the 3D torque-controlled bipedal robot Cassie [116]. Subsequent work explored two principal learning paradigms: reference-based learning, which leverages TO-generated data or motion-capture data to guide policy training [107, 19, 104, 126]; and reference-free learning, where policies are trained entirely from scratch to autonomously discover control strategies [98]. These developments demonstrate that end-to-end frameworks can achieve robust and versatile locomotion skills across complex terrains and dynamic environments [25, 105, 63].
@@
Similarly, hierarchical structures have garnered significant interest. Within this subset, the hybrid approach combines RL-based and model-based methods to enhance both planning and control strategies. Hybrid architectures often integrate learning-based and model-based modules to combine adaptability with physical consistency. One representative design couples a learned HL planner with an LL model-based controller, forming a cascade-structure or deep-planning hybrid scheme [60, 23, 17]. Alternatively, DRL feedback-control hybrids embed learned control policies within model-based feedback loops to enhance tracking precision and disturbance rejection [101, 112]. Learned hierarchical control schemes [3] decompose locomotion into multiple layers, each focusing on specific functions such as navigation and fundamental locomotion skills [84, 83, 127]. To provide a clearer overview of the current landscape, we categorise existing DRL frameworks as shown in Fig. 2.
@@
Current progress across both end-to-end and hierarchical paradigms indicates that a unified framework for DRL-based bipedal locomotion is still far from being realised. Establishing such a framework is essential for consolidating diverse learning pipelines, standardising evaluation metrics, and enabling transferable locomotion capabilities across different robot morphologies. As locomotion tasks become increasingly complex, ranging from basic stabilisation to dynamic parkour and loco-manipulation, the need for consistent benchmarking has intensified.The DARPA Robotics Challenge exemplified this trend by introducing one of the first large-scale evaluation platforms for bipedal humanoids performing real-world locomotion and manipulation tasks [4], highlighting the importance of robustness and practical deployment.
Although several reviews discuss RL for general robotics [54] and model-based methods for bipedal robots [37, 91, 14], none specifically focus on DRL-based frameworks for bipeds. To address this gap, this survey reviews relevant literature according to the following selection criteria: (1) studies that investigate DRL frameworks specifically for bipedal robots; (2) research involving both simulated and physical bipedal robots; (3) approaches that improve policy transfer from simulation to real-world environments; and (4) publications from the last five years (2018–April 2024) sourced from reputable databases and conferences, including Google Scholar, IEEE Xplore, Web of Science, arXiv, and major robotics venues such as CoRL, RSS, ICRA, IROS, and Humanoids.
@@ II End-to-end framework
The end-to-end DRL framework represents a holistic approach in which a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly maps sensory inputs $\mathcal{X}$, such as images, LiDAR data, or proprioceptive feedback [86], together with user commands [98] or pre-defined references [61], into joint-level control actions $\mathcal{U}$. Here, $\mathcal{X}$ represents the sensory input space, $\mathcal{U}$ refers to the space of control actions, and $\pi(\cdot)$ denotes the policy function. This framework obviates the need for manually decomposing the problem into sub-tasks, streamlining the control process.
@@
Methods Works Capabilities Characteristic Advantages and Disadvantages Implementation Flow Chart
Residual learning [115] [116] Forward walk unidirectional walk Adding a residual term to the known motor positions at the current time step. A: Fast convergence speed D: Requires high-quality predefined reference, limits to specific motions, and lacks robustness to complex terrains. [Uncaptioned image]
Guided learning [100] [61] [62] [63] Forward walk Versatile walk Versatile jump Versatile motions Mimic the predefined reference and directly specifies joint-level commands. A: Accelerates the learning process and is robust across terrains. D: Limits to the predefined motions and lacks adaptability to unforeseen changes in environment. [Uncaptioned image]
Reference-free learning [98] [99] [25] Periodic motions Stepping stones walk Visual walk Learn locomotion skills from scratch without any prior knowledge. A: High potential for gait exploration, highly robust to complex terrain D: Requires intensive reward shaping for gait patterns and is relatively expensive in computational resources [Uncaptioned image]
@@ II-A1 Residual learning
Introduced in 2018, a residual learning framework for the bipedal robot Cassie marked a significant advancement [115]. This framework allowed the robot to walk forward by incorporating a policy trained via Proximal Policy Optimisation (PPO) algorithms, as detailed in Appendix A. The policy receives the robot’s states and reference inputs, outputting a residual term that augments the reference at the current timestep. These modified references are then processed by a PD controller to set the desired joint positions. Although this framework has improved the robot’s ability to perform tasks beyond standing [119], it has yet to be physically deployed on a bipedal robot. As a result, it remains impractical for managing walking at varying speeds and is constrained to movement in a single direction.
To transition this framework to a real robot, a sim-to-real strategy based on the previous model was demonstrated, where the policy, trained through a residual learning approach, was subsequently applied to a physical bipedal robot [116]. Compared to model-based methods, this training policy achieves faster running speeds on the same platform, underlining the considerable potential of DRL-based frameworks. However, the robot’s movements remain constrained to merely walking forward or backward.
A unique residual learning approach was introduced to enable omnidirectional walking, where the policy adds a residual term to the current joint positions, allowing gradual omnidirectional walking [92]. In this case, the desired reference is the robot’s current joint positions, which makes the approach distinctive. However, this also limits the policy’s ability to explore more diverse motions, restricting it to a single slow walking pattern.
Residual learning enhances an existing control policy by taking current joint positions or states and applying a residual action to adjust reference actions for better performance. Compared to other learning approaches that directly output joint positions, it is highly sample efficient [23]. However, when predefined references are unstable or of low quality, residual learning may struggle, especially on complex terrains, as the action space is bounded by the reference, limiting the ability to handle unpredictable or uneven terrains.
@@ II-A2 Guided learning
A sim-to-real framework that employs periodic references to initiate the training phase was proposed in [100]. In this framework, the action space directly maps to the joint angles, and desired joint positions are managed by joint PD controllers. The framework also incorporates a Long Short-Term Memory (LSTM) network, as detailed in the Appendix A, which is synchronised with periodic time inputs. However, this model is limited to a single locomotion goal: forward walking. A more diverse and robust walking DRL framework that includes a Hybrid Zero Dynamics (HZD) gait library was demonstrated [61], achieving a significant advancement by enabling a single end-to-end policy to facilitate walking, turning, and squatting.
@@
Despite these advancements, the parameterisation of reference motions introduces constraints that limit the flexibility of the learning process and the policy’s response to disturbances. To broaden the capabilities of guided learning policies, a framework capable of handling multiple targets, including jumping, was developed [62]. This approach introduced a novel policy structure that integrates long-term input/output (I/O) encoding, complemented by a multi-stage training methodology that enables the execution of complex jumping manoeuvrers. An adversarial motion priors approach, employing a style reward mechanism, was also introduced to facilitate the acquisition of user-specified gait behaviours [126]. This method improves the training of high-dimensional simulated agents by replacing complex hand-designed reward functions with more intuitive controls.
@@
While previous works primarily focused on specific locomotion skills, a unified framework that accommodates both periodic and non-periodic motions was further developed [63] based on the foundational work in [62]. This framework enhances the learning process by incorporating a wide range of locomotion skills and introducing a dual I/O history approach, marking a significant breakthrough in creating a robust, versatile, and dynamic end-to-end framework. However, experimental results indicate that the precision of locomotion features, such as velocity tracking, remains suboptimal.
@@
Guided learning methods expedite the learning process by leveraging expert knowledge and demonstrating the capacity to achieve versatile and robust locomotion skills. Through the comprehensive evaluation [63], it is demonstrated that guided learning employs references without complete dependence on them. Conversely, residual learning exhibits failures or severe deviations when predicated on references of inferior quality. This shortfall stems from the framework’s dependency on adhering closely to the provided references, which narrows its learning capabilities.
@@
However, the benefits of reference-based learning come with inherent limitations. Reliance on predefined trajectories often confines the policy to specific gaits, restricting its capacity to explore a broader range of motion possibilities [61, 111]. Moreover, such methods exhibit reduced adaptability when confronted with novel environments or unforeseen perturbations. These limitations are further compounded by the difficulty of acquiring high-quality and task-relevant demonstrations.
@@
Common sources of prior knowledge include TO [60, 36, 61, 63], human motion capture [19], teleoperation [97, 31], and scripted controllers [83]. While informative, these demonstrations often require adaptation due to embodiment mismatch or limited generalisability. Motion retargeting [82, 104, 6], as one of promising direction, addresses this by converting human-centric motions into robot-feasible trajectories, yet it still struggles with preserving fidelity and adapting across morphologies.